# Finer: Investigating and Enhancing Fine-Grained Visual Concept Recognition in Large Vision Language Models

### v2

@@ A.4 Construction of the Instruction-Tuning Mixture
To evaluate the validity of our proposed benchmark, Finer, we construct an instruction-tuning mixture on top of LLaVA-1.5’s instruction-tuning mixture. We use the iNaturalist-2021 dataset to construct the mixture, sampling 20,000 instances from each of the fine-grained instances and their attributes into the training data. We structure each instruction-tuning instance in the AttrSeek pipeline format, with each instruction consistin of three turns: (i) Asking the model for the superordinate concept category, (ii) asking the model to generate a set of external, descriptive visual attributes of the superordinate concept, and finally (iii) predicting the fine-grained concept category given the superordinate concept and the generated attributes set. For each of the three steps, we use GPT-4V to generate 15 possible paraphrases of the instruction in order to avoid biasing the model to specific textual instructions and to retain the model’s instruction-following ability. We trained the model for 1 epoch, which consists of 12k steps; the checkpoint at 12k steps is the one we used to evaluate the fine-grained classification performance in Table 5.

### v4

@@ A.4 Construction of the Instruction-Tuning Mixture
To evaluate the validity of our proposed benchmark, Finer, we construct six different instruction-tuning mixtures on top of LLaVA-1.5’s instruction-tuning mixture. For example, to build a training mixture to train the model for iNaturalist evaluation, i.e., the Finer’s iNaturalist subset, we hold out the iNaturalist dataset for evaluation and include the rest of the five other FGVC datasetes into the instruction-tuning mixture. We use all the six FGVC datasets to construct the mixture, sampling 2.5K instances from each of the datasets and their attributes into the training data; note, the 2.5k instances are sampled to follow a uniform distribution for the number of classes for each dataset. We structure each instruction-tuning instance into the AttrSeek pipeline format, with each instruction consisting of three turns: (i) Asking the model for the coarse-level concept category given the superordinate concept, “Can you identify the bird shown in this image?"; (ii) asking the model to generate a set of external, descriptive visual attributes of the coarse-level concept, “What kind of external descriptive attributes do you see from the penguin", and finally (iii) predicting the fine-grained concept category given the coarse-level concept and the self-generated attributes set. For each of the three steps, we use GPT-4V to generate 15 possible paraphrases of the instruction in order to avoid biasing the model to specific textual instructions and to retain the model’s instruction-following ability. We trained the models for 1 epoch each; the checkpoint for 1 epoch is the one we used to evaluate the FGVC performance in Table 4.
@@ B.2 Knowledge Probing Prompts
Dataset Name Total # of Instances (Test / Training) # of Superordinate Categories # of Coarse Categories # of Fine Categories # of Images Annotation Granularity Hierarchy Partial Attribute Annotation iNaturalist-2021 100K / 2.6M 11 1,103 10K 3,286,843 × CUB-200-2011 5,794 / 5,994 1 59 200 11,788 × FGVC-Aircraft 3,333 / 6,667 1 30 100 10,000 × Stanford Dogs 8,580 / 12,000 1 - 120 20,580 × × NABirds 24,633 / 2,929 1 146 404 48,562 × Stanford Cars 8,144 / 8,041 1 7 196 16,185 × × FINER 372K/2.63M 16 1,416 11,171 3,393,958
@@ B.4 Coarse-Grained Label Generation Prompts
B.5 Linear Probing Experiment
To evaluate the quality of output representations before and after the multimodal projection in LLaVA-1.5 (7B), we perform linear probing over the output representations of the vision encoder (CLIP-ViT-L/14 Radford et al. (2021)) and the projected representations in the textual space. The experimental settings are 10 epochs for finetuning and we use accuracy as the evaluation metric. We train on 4x V100 16GB for 57 GPU hours.